suppose ha he money demand function is (M/P)d =1000-100r, where r is he interest rate in percent. the money supply M is 1000 and the price level P is N$2.
a) graph the supply and demand for real money balances.
b)what is the equilibrium interest rate?
c) assume that the price level is fixed. what happens to the equilibrium interest rate if the supply of money is raised by 39%?
d) if the Fed wishes to raise the interest rate to 7%, what money supply should it set?
1
Expert's answer
2015-09-19T00:00:43-0400
(M/P)d =1000-100r, Ms = 1000, P = N$2.
a) https://encrypted-tbn3.gstatic.com/images?q=tbn:ANd9GcTdTnoNRs57uqJVlKj8AZfS_g2RJ8Mf3AX3uhMKsNRPYa8pI0a3Uw
b) the equilibrium interest rate will be:
Md = Ms, so:
2*(1000 - 100r) = 1000
r = 5%
c) If the price level is fixed and the supply of money is raised by 39%, then the equilibrium interest rate will be:
2*(1000 - 100r) = 1390
r = 3.45%
d) if the Fed wishes to raise the interest rate to 7%, the money supply should be:
2*(1000 - 7*100) = 600.